OpenStudy (anonymous):

PLEASE HELP ME!!!!!! Eliana is reading a 920-page book. She reads 20 pages per day in the book. If x represents the number of days Eliana has spent reading the book so far, which of the following equations can be used to find the number of pages remaining in the book for Eliana to read? A. y = 20 - 920x B. y = 920 - 20x C. y = 20 + 920x D. y = 920 + 20x @student9626 @studygurl14

OpenStudy (anonymous):

b

OpenStudy (anonymous):

20 pgs per day, and 920 total, right?

OpenStudy (lifeisadangerousgame):

Ooo, I love these! x represents that number of days she has spent reading so far. She reads 20 pages per day and has 920 pages to read altogether. 20 is our independent variable, which means that it changes depending on what number x is. So y = 920 - 20x.

OpenStudy (anonymous):

Correct! Thank you!!!

OpenStudy (lifeisadangerousgame):

@SickTwistd , we are not supposed to give them the answer but help them get to the answer. If we just give them it, they won't learn :/

OpenStudy (anonymous):

so lets say 7 days, 7*20=140, 920-140=780, which is the remaining number of pages

OpenStudy (anonymous):

CANT YOU SEE THA XPLANATION? @lifeisadangerousgame

OpenStudy (anonymous):

lol

OpenStudy (lifeisadangerousgame):

The "xplanation" you're giving only tells them what 7*20 is. You have to show them how the problem works out.

OpenStudy (lifeisadangerousgame):

But the OP is happy and has their answer, so see ya